# JL1-CD: A New Benchmark for Remote Sensing Change Detection and a Robust Multi-Teacher Knowledge Distillation Framework

### v2

@@ Abstract
Deep learning has achieved significant success in the field of remote sensing image change detection (CD), yet two major challenges remain: the scarcity of sub-meter, comprehensive open-source CD datasets, and the difficulty of achieving consistent and satisfactory detection results across images with varying change areas. To address these issues, we introduce the JL1-CD dataset, which consists of 5,000 pairs of 512 × 512 pixel images with a resolution of 0.5 to 0.75 meters. This all-inclusive dataset covers a wide range of human-induced and natural changes, including buildings, roads, hardened surfaces, woodlands, grasslands, croplands, water bodies, and photovoltaic panels, among others. Additionally, we propose a novel multi-teacher knowledge distillation (MTKD) framework that leverages the Origin-Partition (O-P) strategy to enhance CD performance. In the O-P strategy, we partition the training data based on the Change Area Ratio (CAR) to train separate models for small, medium, and large CAR values, alleviating the learning burden on each model and improving their performance within their respective partitions. Building upon this, our MTKD framework distills knowledge from multiple teacher models trained on different CAR partitions into a single student model,enabling the student model to achieve superior detection results across diverse CAR scenarios without incurring additional computational or time overhead during the inference phase. Experimental results on the JL1-CD and SYSU-CD datasets demonstrate that the MTKD framework significantly improves the performance of CD models with various network architectures and parameter sizes, achieving new state-of-the-art results. The JL1-CD dataset and code are available at https://github.com/circleLZY/MTKD-CD.
Index Terms:
Knowledge distillation, change detection, remote sensing.
I INTRODUCTION
@@
Early traditional CD methods primarily relied on image processing techniques, detecting changes by directly comparing pixel values or spectral features between multi-temporal images. Examples include Image Differencing[2], Change Vector Analysis (CVA)[3], Principal Component Analysis (PCA)[4], Kauth–Thomas (KT) transforms[5], and Multivariate Alteration Detection (MAD)[6]. While these methods are simple and intuitive, they exhibit limited performance when handling complex change patterns, such as those affected by significant noise, lighting variations, or seasonal differences.
With the rise of machine learning (ML), CD methods began to incorporate feature extraction and classifiers. Common techniques include Support Vector Machines (SVM)[7], Random Forest (RF)[8], K-means clustering[9] and so on. These machine learning approaches significantly improved detection accuracy but heavily relied on high-quality labeled data and manually designed features.
In recent years, the rapid advancement of deep learning (DL) has revolutionized remote sensing CD, delivering substantial performance breakthroughs. DL-based CD methods generally involve three steps: (1) extracting change features from image pairs, (2) generating change maps based on the extracted features, and (3) predicting labels based on the feature maps. Convolutional Neural Networks (CNNs), which achieved remarkable success in image processing, were the first neural network architecture applied to remote sensing CD and remain widely optimized and utilized today [10, 11, 12]. With the introduction of Transformers, some studies have explored their application in CD tasks [13, 14]. More recently, Foundational Model (FM) has emerged as a novel paradigm, aiming to achieve multi-task and multi-domain generalization through large-scale pretraining [15].
However, DL-based CD methods generally face two major challenges: the scarcity of high-quality, high-resolution, all-inclusive CD datasets and limitations in handling highly dynamic change areas. Although numerous CD datasets have been constructed and proposed, they are often tailored to specific scenarios, which restricts the generalization capabilities of the algorithms. For instance, models trained on datasets focused on human-induced changes often fail to perform effectively when confronted with natural change scenarios. On the other hand, the learning capacity of these models is inherently limited. Most existing algorithms rely on a single-phase training approach, typically end-to-end training. While such training strategies can achieve satisfactory results within a constrained range of changes, the models’ performance significantly degrades when addressing scenarios with wide variations in change areas.
@@
To address the aforementioned challenges, we construct a new large-scale, high-resolution, all-inclusive open-source CD dataset, named “JL1-CD” (named after the Jilin-1 satellite). This dataset comprises 5,000 pairs of 512 × 512 images captured in China, with a resolution of 0.5–0.75 meters, along with binary change labels at the pixel level. The JL1-CD dataset not only includes common human-induced changes such as buildings and roads but also encompasses various natural changes, such as forests, water bodies, and grasslands. Additionally, we propose a multi-teacher knowledge distillation (MTKD) framework for change detection optimization. First, we introduce the O-P strategy. To address the difficulty of handling highly dynamic change areas in existing algorithms, we propose the concept of the Change Area Ratio (CAR) and partition the dataset based on different CAR levels. CD models are then trained on each partition, reducing the learning burden on individual models, thereby achieving better training outcomes and higher detection accuracy. Next, to lower the computational and time complexity during inference, we extend the O-P strategy by training a student model under the MTKD framework. The student model learns the strengths of teacher models optimized for various CAR scenarios, achieving superior detection accuracy without increasing resource consumption during inference.
@@
  1. 1)
    We introduce JL1-CD, a new sub-meter, all-inclusive open-source CD dataset comprising 5,000 pairs of remote sensing image patches with a resolution of 0.5–0.75 meters.
  2. 2)
    We propose the O-P strategy, which partitions the training of CD models based on CAR levels, significantly improving performance across diverse CAR scenarios.
  3. 3)
    We further develop the MTKD framework, where models trained under the O-P strategy serve as teacher models. The student model trained under the supervision of multiple teachers achieves superior detection accuracy without additional computational or time costs during inference.
  4. 4)
    Extensive benchmarking experiments on existing algorithms demonstrate that O-P and MTKD significantly enhance performance across various architectures and parameter sizes, achieving new state-of-the-art (SOTA) results.
II RELATED WORKS
II-A Traditional and ML-Based CD
Traditional CD methods have been extensively studied in remote sensing, with early approaches relying on simple algebraic operations such as image differencing [2] and image ratioing [16]. These techniques compute pixel-level differences or ratios between two images and apply a threshold to identify change regions. Subsequent advancements introduced improved thresholding strategies, such as the Otsu method [17] and the normalized difference vegetation index (NDVI) algorithm [18], to enhance detection accuracy. Transform-based methods, such as PCA [4] and MAD [6], were later adopted, leveraging statistical properties of images. However, these methods are heavily dependent on image statistics, which limits their scalability and precision in large-scale, high-resolution CD applications. The advent of machine learning has significantly enhanced the ability to extract useful change features. For instance, Bovolo et al. [19] proposed an unsupervised change detection framework that leverages a semi-supervised SVM initialized with pseudo-training data, effectively addressing the complexity of multi-temporal spectral feature analysis. Wessels et al. [20] developed an automated system for land cover update mapping, integrating iteratively reweighted MAD (IRMAD) for change mask generation and RF classifiers for robust land cover classification, achieving notable accuracy in operational settings. Despite these advancements, traditional and early ML-based approaches often rely on manually designed features, which perform well in straightforward scenarios but exhibit limited generalization capability for complex and diverse change types.
II-B DL-Based CD
@@
In recent years, deep learning has experienced rapid advancements, achieving remarkable success in remote sensing image CD. As illustrated in Fig. 1, we present a timeline of the development of mainstream DL-based CD algorithms. Based on the differences in neural network architectures and training paradigms, DL-based CD methods can be classified into three main categories
CNN-Based CD
CNNs serve as the foundation of many early DL-based CD methods and remain widely utilized today. Daudt et al. [10] proposed three fully convolutional neural network architectures, including two Siamese network extensions, which achieved significant improvements in accuracy and efficiency for CD tasks on multiple datasets. Zhang et al. [21] introduced the Image Fusion Network (IFN), which employs a deeply supervised two-stream architecture for high-resolution remote sensing CD, achieving SOTA performance with superior boundary completeness and compactness in change maps. Chen et al. [22] proposed the Siamese-based Spatial-Temporal Attention Network (STANet), incorporating a novel CD self-attention module to model spatial-temporal dependencies at various scales, significantly improving F1-scores on benchmark datasets. Fang et al. [23] designed SNUNet-CD, a densely connected Siamese network that preserves localization information and employs an Ensemble Channel Attention Module (ECAM) for deep supervision, achieving better trade-offs between accuracy and computational cost. Zheng et al. [24] proposed ChangeStar, a model leveraging single-temporal supervised learning with ChangeMixin modules to train CD models using unpaired images. Han et al. [25] introduced HANet, a hierarchical attention network with progressive foreground-balanced sampling and a lightweight self-attention mechanism, effectively addressing class imbalance in CD tasks and achieving superior results on highly imbalanced datasets. The Change Guiding Network (CGNet) introduced by Han et al. [11] utilizes a self-attention mechanism to improve edge detection and internal consistency in change maps, demonstrating robust performance across multiple CD datasets. Some studies have focused on designing lightweight and fast CD models. Codegoni et al. [26] presented TinyCD, a lightweight and efficient CD model using a Siamese U-Net architecture and the Mix and Attention Mask Block (MAMB), outperforming SOTA models while being significantly smaller and faster. Xing et al. [27] proposed LightCDNet, a lightweight CD model with an early fusion backbone and pyramid decoder.
Figure 2: Sample images from the JL1-CD dataset. Each row, from top to bottom, represents: the image at time 1, the image at time 2, and the ground truth label. Each column corresponds to different change types: (a) Decrease in woodland; (b) Building changes; (c) Conversion of cropland to greenhouses; (d) Road changes; (e) Waterbody changes; (f) Surface hardening; and (g) Photovoltaic panel construction
Transformer-Based CD
@@
Transformer-based methods have emerged as a promising approach for CD. Chen et al. [28] introduced the bitemporal image transformer (BIT), combining a transformer encoder with a ResNet backbone to model spatial-temporal contexts efficiently. Bandara et al. [13] proposed ChangeFormer, a fully transformer-based Siamese network for CD, which unifies a hierarchical transformer encoder with a multi-layer perceptron (MLP) decoder. Fang et al. [14] introduced the Changer series framework, a novel architecture for CD that incorporates alternative interaction layers between bi-temporal features. This framework is applicable to both CNN-based and Transformer-based models, significantly enhancing the performance of the original models.
FM-Based CD
@@
Recently, foundation models have become a new training paradigm. There have been works utilizing remote sensing data to fine-tune pretrained models such as Vision Transformers (ViT)[29], Segment Anything Model (SAM)[30], and Contrastive Language-Image Pretraining (CLIP)[31], achieving higher performance in CD tasks[32, 33, 34, 35]. Li et al. [32] proposed the Bi-Temporal Adapter Network (BAN), a universal FM-based framework for CD, which enhances existing models with minimal additional parameters and achieves significant performance improvements. Chen et al. [34] introduced Time Travelling Pixels (TTP), a method that integrates latent knowledge from the SAM model into CD, overcoming domain shifts and spatio-temporal complexities, demonstrating SOTA results on the LEVIR-CD[22] dataset. Zheng et al. [35] developed AnyChange, a zero-shot CD model built on the SAM that utilizes bitemporal latent matching for training-free adaptation, setting a new SOTA on the SECOND[36] benchmark and achieving significant improvements in both unsupervised and supervised CD tasks.
II-C Knowledge Distillation in CD
@@
Yan et al. [40] proposed a novel self-supervised learning approach for unsupervised CD by fusing domain knowledge of remote sensing indices during both training and inference. By calculating cosine similarity, they selected high-similarity feature vectors from both the teacher and student networks to implement a hard negative sampling strategy, effectively improving CD performance. Wang et al. [41] addressed remote sensing semantic CD (SCD), which focuses on detecting changes in land cover and land use over time. The authors introduced a dual-dimension feature interaction network (DFINet) that enhances intraclass and interclass feature differentiation by incorporating a temporal difference feature enhancement (TDFE) module, which captures temporal features comprehensively. Wang et al. [42] proposed a KD-based method for CD (CDKD), designed to overcome the challenges of deploying large deep learning models with high computational and storage requirements on resource-constrained spaceborne edge devices. Although these methods have successfully utilized KD to enhance the performance of various student models, they are tailored to specific models and do not provide a generalized distillation framework applicable to various CD models. Furthermore, there is a lack of open-source KD-based code for remote sensing image CD tasks.
In contrast, the proposed MTKD framework significantly improves the performance of CD models with various architectures and parameter sizes, and we commit to open-sourcing all the code and models.
TABLE I: Information of Open-Source CD Datasets and the Proposed JL1-CD Dataset
Dataset Class Image Pairs Image Size Resolution
SZTAKI[43] 1 13
@@
DSIFN[21] 1 394 512 × 512 2
SECOND[36] 6 4,662 512 × 512 0.5-3
WHU-CD[44] 1 1 32,20 × 15,354 0.2
LEVIR-CD[22] 1 637 1,024 × 1,024 0.3
S2Looking[45] 1 5,000 1,024 × 1,024 0.5-0.8
CDD[46] 1 16,000 256 × 256 0.03-1
SYSU-CD[47] 1 20,000 256 × 256 0.5
JL1-CD 1 5,000 512 × 512 0.5-0.75
@@
III JL1-CD Dataset
We collect the number of types of changes, number of image pairs, image size, and resolution information of mainstream CD datasets in Table I. The SZTAKI AirChange Benchmark [43] contains 12 pairs of 952 × 640 and one pair of 1,048 × 724 optical aerial images. It is one of the earliest and most commonly used CD datasets in early research. The DSIFN dataset [21] consists of 6 large bi-temporal image pairs from 6 cities in China, which are cropped into 394 sub-image pairs, each sized 512 × 512. The SECOND dataset [36] includes 4,662 pairs of aerial images collected from multiple platforms and sensors, covering cities such as Hangzhou, Chengdu, and Shanghai. Unlike other datasets that classify changes into only two categories (change and no change), SECOND provides detailed annotations for change types, including non-vegetated surfaces, trees, low vegetation, water, buildings, and playgrounds. However, the resolution of all or part of the images in these datasets does not reach the sub-meter level. WHU-CD [44], LEVIR-CD [22], and S2Looking [45] are very popular datasets that are specifically designed for monitoring building changes. These datasets predominantly include human-induced changes and lack natural change types. The CDD dataset[46] is derived from 7 pairs of 4,725 × 2,700 real-world seasonal change remote sensing images. The SYSU-CD dataset[47] contains 20,000 pairs of 0.5-meter aerial images captured in Hong Kong between 2007 and 2014. These datasets feature high resolution and diverse change types.
To provide a better benchmark for evaluating CD algorithms, we propose the JL1-CD dataset, a high-resolution, all-inclusive change detection dataset. JL1-CD includes 5,000 pairs of satellite images captured in China from early 2022 to the end of 2023, including Shandong, Ningxia, Anhui, Hebei, Hunan, and other regions. The images have sub-meter resolutions ranging from 0.5 to 0.75 meters and are sized 512 × 512 pixels. As shown in Fig. 2, the dataset covers various common human-induced and natural surface features, such as buildings, roads, hardened surfaces, woodlands, grasslands, croplands, water bodies and photovoltaic (PV) panels. It is worth noting that land cover changes induced by the construction of PV panels represent a significant change type in this dataset. As illustrated in Fig. 2 (g), these panels appear as black rectangles in the imagery due to their efficient sunlight absorption design. The resulting change regions exhibit diverse shapes and sizes, effectively enriching the diversity of change samples in our dataset. The original 5,000 image pairs are divided into 4,000 pairs for training and 1,000 pairs for testing, following an 80:20 split. The JL1-CD dataset will be made openly available for all research needs.
Figure 4: Sample images with different change area ratios (CAR). Each column represents a specific CAR: (a) 0.00%; (b) 19.98%; (c) 39.93%; (d) 59.96%; (e) 80.25%; and (f) 100.00%.
Figure 5: CAR distribution of the training, validation and test sets in JL1-CD.
Figure 6: CAR distribution of the training and test sets in SYSU-CD.
IV Methodology
In this section, we provide a comprehensive overview of the proposed methods. In Section IV-A, we first introduce the Origin-Partition (O-P) strategy designed for the challenging all-inclusive CD dataset. Building upon the O-P strategy, we further present our Multi-Teacher Knowledge Distillation (MTKD) framework in Section IV-B. Finally, in Section IV-C, we describe the overall loss function used for training the teacher and student models.
IV-A O-P Strategy

### v3

@@ Abstract
Change detection (CD) in remote sensing images plays a vital role in Earth observation. However, the scarcity of high-resolution, comprehensive open-source datasets and the difficulty in achieving robust performance across varying change types remain major challenges. To address these issues, we introduce JL1-CD, a large-scale, sub-meter CD dataset consisting of 5,000 image pairs. We further propose a novel Origin-Partition (O-P) strategy and integrate it into a Multi-Teacher Knowledge Distillation (MTKD) framework to enhance CD performance. The O-P strategy partitions the training set by Change Area Ratio (CAR) and trains specialized teacher models on each subset. The MTKD framework then distills complementary knowledge from these teachers into a single student model, enabling improved detection results across diverse CAR scenarios without additional inference cost. Our MTKD approach demonstrated strong performance in the 2024 “Jilin-1” Cup challenge, ranking first in the preliminary and second in the final rounds. Extensive experiments on the JL1-CD and SYSU-CD datasets show that the MTKD framework consistently improves the performance of CD models with various network architectures and parameter sizes, establishing new state-of-the-art results. Code and dataset are available at https://anonymous.4open.science/r/MTKD-A-84B8.
1 Introduction
@@
In recent years, the rapid advancement of deep learning (DL) has revolutionized remote sensing CD, delivering substantial performance breakthroughs. Convolutional Neural Networks (CNNs), which have demonstrated remarkable success in image processing, were the first neural network architecture applied to remote sensing CD and remain widely used and optimized today [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. With the introduction of Transformers, several studies have explored their application in CD tasks [12, 13, 14, 15]. The Mamba architecture, a state space sequence model designed for efficient long-range dependency modeling, has attracted growing interest as a promising alternative to Transformers in CD tasks [16, 17]. More recently, Foundation Model (FM) has emerged as a novel paradigm, aiming to achieve multi-task and multi-domain generalization through large-scale pretraining [18, 19, 20, 21, 22].
However, DL-based CD methods generally face two major challenges: the scarcity of high-quality, high-resolution, all-inclusive CD datasets and limitations in handling highly diverse change scenarios. Although numerous CD datasets have been proposed, they are often tailored to specific scenarios, thus restricting the generalization capabilities of the algorithms. For instance, models trained on datasets focused on human-induced changes often fail to perform effectively when confronted with natural change scenarios. On the other hand, the learning capacity of these models is inherently limited. Most existing algorithms rely on a single-phase training approach, typically end-to-end training. Although such training strategies perform well on limited change types, model accuracy degrades significantly as change diversity increases.
@@
To address the aforementioned challenges, we construct a new large-scale, high-resolution, all-inclusive open-source CD dataset, JL1-CD, which is named after the Jilin-1 satellite. This dataset comprises 5,000 pairs of 512 × 512 images captured in China, with a resolution of 0.5–0.75 meters, along with binary change labels at the pixel level. The JL1-CD dataset not only includes common human-induced changes such as buildings and roads but also encompasses various natural changes, such as forests, water bodies, and grasslands. To quantitatively characterize the diversity of change areas in datasets, we propose the Change Area Ratio (CAR) metric and partition the dataset into subsets based on different CAR levels, which we refer to as the Original-Partition (O-P) strategy. By training separate CD models on each partition, the O-P strategy effectively simplifies the learning task and improves detection accuracy. Building upon this, we further propose the Multi-Teacher Knowledge Distillation (MTKD) framework that distills complementary knowledge from these partition-specific teacher models into a single student model. This student model combines the strengths of individual teachers, achieving superior detection performance without increasing resource consumption during inference.
@@
  • We introduce JL1-CD, a new sub-meter, all-inclusive open-source CD dataset.
  • We evaluate numerous existing algorithms on the JL1-CD dataset, establishing a new benchmark for remote sensing CD by publicly releasing the associated dataset and code. This contribution is anticipated to facilitate further advancements in the field.
  • We propose MTKD, a general optimization framework for CD models, which achieves consistent improvement across diverse CD models with varying architectures and parameter scales on different datasets.
2 Related Works
2.1 DL-Based CD
@@
Deep learning has experienced rapid advancements, achieving remarkable success in remote sensing image CD. As illustrated in Figure 1, we present a timeline of the development of mainstream DL-based CD algorithms. Based on the differences in neural network architectures and training paradigms, we briefly review three major categories of DL-based CD methods that are widely adopted in current research.
CNN-Based CD: CNNs serve as the foundation of many early DL-based CD methods and remain widely used today. Daudt et al. [2] proposed three fully convolutional neural network architectures, marking the entry of DL into CD research. This breakthrough spurred numerous subsequent CNN-based solutions. More recently, researchers have also developed efficient CD models like TinyCD [7] and LightCDNet [10] that achieve superior performance while being significantly smaller and faster.
@@
Transformer-Based CD: Transformer-based methods have emerged as a promising approach for CD. Chen et al. [23] introduced the bi-temporal image transformer (BIT), combining a transformer encoder with a ResNet backbone to model spatial-temporal contexts efficiently. Bandara et al. [12] proposed ChangeFormer, a fully transformer-based Siamese network for CD, which unifies a hierarchical transformer encoder with a multi-layer perceptron (MLP) decoder. Fang et al. [13] introduced the Changer series framework, a novel architecture for CD that incorporates alternative interaction layers between bi-temporal features. This framework is applicable to both CNN-based and Transformer-based models, enhancing the performance of the original models.
@@
FM-Based CD: Recently, foundation models have become a new training paradigm. There have been works utilizing remote sensing data to fine-tune pretrained models such as Vision Transformers (ViT) [24], Segment Anything Model (SAM) [25], and Contrastive Language-Image Pretraining (CLIP) [26], achieving better performance in CD tasks [20, 18, 21]. Li et al. [20] proposed the Bi-Temporal Adapter Network (BAN), a universal FM-based framework for CD, which enhances existing models with minimal additional parameters and achieves significant performance improvements. Chen et al. [21] introduced Time Travelling Pixels (TTP), a method that integrates latent knowledge from the SAM model into CD, overcoming domain shifts and spatio-temporal complexities.
2.2 Knowledge Distillation in CD
@@
Yan et al. [30] proposed a novel self-supervised learning approach for unsupervised CD by fusing domain knowledge of remote sensing indices during both training and inference. Wang et al. [31] addressed remote sensing semantic CD (SCD), which focuses on detecting changes in land cover and land use over time. Wang et al. [32] proposed a KD-based method for CD (CDKD), designed to overcome the challenges of deploying large deep learning models with high computational and storage requirements on resource-constrained spaceborne edge devices. Although these methods have successfully utilized KD to enhance the performance of various student models, they are tailored to specific models and do not provide a generalized distillation framework applicable to other CD models. Furthermore, open-source KD-based implementations for remote sensing CD tasks remain limited.
In contrast, the proposed MTKD framework substantially improves the performance of CD models with various architectures and parameter sizes, and we commit to open-sourcing all the code and models.
3 JL1-CD Dataset
Figure 2: Representative samples from JL1-CD dataset. Each column, from top to bottom, represents the pre-event image, the post-event image, and the ground truth label. Each column corresponds to different change types: (a) decrease in woodland; (b) building changes; (c) conversion of cropland to greenhouses; (d) road changes; (e) waterbody changes; (f) surface hardening; and (g) photovoltaic panel construction.
Table 1: Information of open-source binary CD datasets and the proposed JL1-CD dataset.
Dataset Year Image Pairs Image Size Resolution
SZTAKI [33] 2009 13
@@ 3 JL1-CD Dataset
CDD [34] 2011 16,000 256 × 256 0.03-1
WHU-CD [35] 2018 1 32,20 × 15,354 0.2
DSIFN [3] 2020 394 512 × 512 2
LEVIR-CD [4] 2020 637 1,024 × 1,024 0.3
S2Looking [36] 2021 5,000 1,024 × 1,024 0.5-0.8
SYSU-CD [37] 2021 20,000 256 × 256 0.5
JL1-CD 2025 5,000 512 × 512 0.5-0.75
Table 1 summarizes mainstream binary CD datasets in terms of release year, number of image pairs, image size, and spatial resolution. The SZTAKI AirChange Benchmark [33] contains 12 pairs of 952 × 640 and one pair of 1,048 × 724 optical aerial images. It is one of the earliest and most widely used CD datasets in early research. The DSIFN dataset [3] consists of 6 large bi-temporal image pairs from 6 cities in China, which are cropped into 394 sub-image pairs, each sized 512 × 512. However, the resolution of all or part of the images in these datasets does not reach the sub-meter level. WHU-CD [35], LEVIR-CD [4], and S2Looking [36] are widely recognized datasets focusing on monitoring building changes. These datasets predominantly include human-induced changes and lack natural change types. The CDD dataset [34] is derived from 7 pairs of 4,725 × 2,700 real-world seasonal change remote sensing images. The SYSU-CD dataset [37] contains 20,000 pairs of 0.5-meter aerial images captured in Hong Kong between 2007 and 2014. While both datasets offer high resolution and diverse change types, the image acquisition dates are relatively outdated.
To provide a new benchmark for evaluating CD algorithms, we propose the JL1-CD dataset, a high-resolution, all-inclusive change detection dataset. JL1-CD includes 5,000 pairs of satellite images captured in China from early 2022 to the end of 2023, including Shandong, Ningxia, Anhui, Hebei, Hunan, and other regions. The images have sub-meter resolutions ranging from 0.5 to 0.75 meters and are sized 512 × 512 pixels. As shown in Figure 2, the dataset covers various common human-induced and natural surface features, such as buildings, roads, hardened surfaces, woodlands, grasslands, croplands, water bodies and photovoltaic (PV) panels. The original 5,000 image pairs are divided into 4,000 pairs for training and 1,000 pairs for testing, following a split of 80%: 20%. The JL1-CD dataset will be made openly available for all research needs.
@@
4 Methodology
4.1 O-P Strategy